Integrate (7x+8)/(2x²+11x+5)dx
Integrate (3x²+18x+3)/(3x²+5x-2)dx
Integrate (6x²)/(x²-4)dx
Integrate (sin inverse x) /(x²-4)dx
Integrate (x²+1)³/(x+2)dx
1
Expert's answer
2020-03-02T09:26:11-0500
  1. "I=\int\frac{7x+8}{2x^2+11x+5}dx"

decompose the denominator into factors

"2x^2+11x+5=2(x+5)(x+0.5)\\
2x^2+11x+5=0\\
D=121-4\cdot2\cdot5=81\\
x_1=\frac{-11-9}{4}=-5\\
x_2=\frac{-11+9}{4}=-0.5"

"I=\int\frac{7x+8}{2(x+5)(x+0.5)}dx=\\
=\frac{1}{2}\int\frac{7x+8}{(x+5)(x+0.5)}dx"

decompose into simple fractions

"\frac{7x+8}{(x+5)(x+0.5)}=\frac{A}{(x+5)}+\frac{B}{(x+0.5)}=\frac{A(x+0.5)+B(x+5)}{(x+5)(x+0.5)}\\
1) x=-5: \\
7\cdot(-5)+8=A(-5+0.5)\\
A=6\\
2) x=-0.5\\
7\cdot(-0.5)+8=B(-0.5+5)\\
B=1"

"I=\frac{1}{2}\int(\frac{6}{x+5}+\frac{1}{x+0.5})dx=\\
=\frac{1}{2}(6ln|x+5|+ln|x+0.5|)+C"


2

"I=\int\frac{3x^2+18x+3}{3x^2+5x-2}dx"

perform the division and decompose the denominator into factors

"\begin{matrix}
 3x^2+18x+3| & 3x^2+5x-2 \\
 3x^2+5x-2 & 1\\
-----\\
13x+5\\
\end{matrix}\\
\frac{3x^2+18x+3 }{3x^2+5x-2} =1+\frac{13x+5}{3x^2+5x-2 }\\
3x^2+5x-2 =3(x+2)(x-\frac{1}{3})\\
3x^2+5x-2 =0\\
D=25+4\cdot3\cdot2=49\\
x_1=\frac{-5-7}{6}=-2\\
x_2=\frac{-5+7}{6}=\frac{1}{3}"

"I=\int(1+\frac{13x+5}{3(x+2)(x-\frac{1}{3})})dx=\\
=x+\frac{1}{3}\int\frac{13x+5}{(x+2)(x-\frac{1}{3})}"

decompose into simple fractions

"\frac{13x+5}{(x+2)(x-\frac{1}{3})}=\frac{A}{(x+2)}+\frac{B}{(x-\frac{1}{3})}=\frac{A(x-\frac{1}{3})+B(x+2)}{(x+2)(x-\frac{1}{3})}\\
1) x=-2: \\
13\cdot(-2)+5=A(-2-\frac{1}{3})\\
A=\frac{33}{7}\\
2) x=\frac{1}{3}\\
13\cdot\frac{1}{3}+5=B(\frac{1}{3}+2)\\
B=4"

"I=x+\frac{1}{3}\int(\frac{\frac{33}{7}}{x+2}+\frac{4}{x-\frac{1}{3}})dx=\\
=x+\frac{1}{3}(\frac{33}{7}ln|x+2|+4ln|x-\frac{1}{3}|)+C"



3

"\int\frac{\sin^{-1}x}{x^2-4}dx=\frac{1}{8}(-2i(Li_2(i(-2+\sqrt{3})e^{-i\sin^{-1}x})+\\
+Li_2(-i(2+\sqrt{3})e^{-i\sin^{-1}x}))+\\
+2i(Li_2(-i(-2+\sqrt{3})e^{-i\sin^{-1}x})+\\
+Li_2(-i(2+\sqrt{3})e^{-i\sin^{-1}x}))+\\
+(-2\sin^{-1}x+\pi+4sin^{-1}\sqrt{\frac{3}{2}})\cdot\\
\cdot \log(1-i(\sqrt{3}-2)e^{-i\sin^{-1}x})+\\
+(-2\sin^{-1}x+\pi+4sin^{-1}\sqrt{\frac{3}{2}})\cdot\\
\cdot \log(1+i(\sqrt{3}+2)e^{-i\sin^{-1}x})+\\
+\log(x-2)(\pi-2\sin^{-1}x)+\\
+2\log(x-2)\sin^{-1}x-\\
-\log(x+2)(\pi-2\sin^{-1}x)-\\
-2\log(x+2)\sin^{-1}x-\\
-\log(1+i(\sqrt{3}-2)e^{-i\sin^{-1}x})\cdot\\
\cdot(-2\sin^{-1}x+\pi-4i\sinh^{-1}\frac{1}{\sqrt{2}})-\\
--\log(1-i(\sqrt{3}+2)e^{-i\sin^{-1}x})\cdot\\
\cdot(-2\sin^{-1}x+\pi+4i\sinh^{-1}\frac{1}{\sqrt{2}})+\\
+8isin^{-1}\sqrt{\frac{3}{2}} \tan^{-1}(\frac{\cot(\frac{1}{4}(2\sin^{-1}x+\pi))}{\sqrt{3}})-\\
- 8\sinh^{-1}\frac{1}{\sqrt{2}} \tan^{-1}(\sqrt{3}\cot(\frac{1}{4}(2\sin^{-1}x+\pi))))+c"

"\sin^{-1}x" is the inverse sine function

"Li_nx" isthe polylogarithm function



4

"\int\frac{6-x}{x^2-4}dx=\int\frac{6}{x^2-4}dx-\int\frac{x}{x^2-4}dx=\\
=6\cdot\frac{1}{4}\ln|\frac{x-2}{x+2}| -\frac{1}{2}\int\frac{d(x^2-4)}{x^2-4}dx=\\
=\frac{3}{2}\ln|\frac{x-2}{x+2}| -\frac{1}{2}\ln|x^2-4|+C"





5

"I=\int\frac{(x^2+1)^3}{x+2}dx=\int\frac{x^6+3x^4+3x^2+1}{x+2}dx"

decompose the numerator by degrees "x+2"

"\begin{matrix}
 & 1&0&3&0&3&0&1 \\
 -2 & 1&-2&7&-14&31&-62&125\\
-2&1&-4&15&-44&119&-300\\
-2&1&-6&27&-98&315\\
-2&1&-8&43&-184\\
-2&1&-10&63\\
-2&1&-12\\
-2&1
\end{matrix}"

then

"(x+2)^6-12(x+2)^5+63(x+2)^4-\\
-184(x+2)^3+315(x+2)^2-300(x+2)+125"

"I=\int\frac{(x+2)^6-12(x+2)^5+63(x+2)^4-184(x+2)^3+315(x+2)^2-300(x+2)+125}{x+2}dx=\\
=\int((x+2)^5-12(x+2)^4+63(x+2)^3-\\
-184(x+2)^2+315(x+2)-300+\frac{125}{x+2})dx=\\
=\frac{(x+2)^6}{6}-12\cdot\frac{(x+2)^5}{5}+63\cdot\frac{(x+2)^4}{4}-\\
-184\cdot\frac{(x+2)^3}{3}+315\cdot\frac{(x+2)^2}{2}-300x+\\
+125ln|x+2|+C"